OpenStudy (anonymous):

I really need help, i am extremely confused, and have only right now that i can complete this so please help

OpenStudy (cwrw238):

I'll give you an example of how you would represent a bed:- say that the bed measures 2 meters by 3 meters and we can represent it as a rectangle then the area = 2*3 = 6 meters now a convenient scale would be 1 meter = 2 cms so your scale drawing would be 3*2 = 6 cms long and 2*2 = 4 cms wide so you would draw a rectangle with these dimensions hope this helps

OpenStudy (cwrw238):

the area of your scale drawing would be 4*6 = 24 sq cms

OpenStudy (cwrw238):

if you want to make a smaller scale model then use 1 meter = 1 cm

OpenStudy (cwrw238):

If you want 6 objects then maybe that smaller scale would be better